Mr. Toriop owns 5,000 shares of stock in Yummy Corporation. The company has announced that it will pay a dividend of $5 per share in one year and a liquidating dividend of $50 per share in two years. The required return on Yummy stock is 12%.   

a. What is the current share price of your stock?   

b. What will be the company’s share price in one year’s time?

c. Mr. Toriop wishes to have equal amounts of dividend income for the next two years. How can he use homemade dividends to achieve this goal? Check that the present value of the cash flows will be the same as it is before the homemade dividends. (Hint: Dividends will be in the form of an annuity.)

d. Suppose Mr. Toriop is thinking about buying a house for $220,000 in one year. How can he use homemade dividends to achieve this goal? Check that the present value of the cash flows will be the same as it is before the homemade dividends.   

e. Suppose Mr. Toriop is thinking about postponing the house purchase for two years, by which time the price of the house will have increased by $46,800. How can he use homemade dividends to achieve this goal? Check that the present value of the cash flows will be the same as it is before the homemade dividends.   
Answer #1
a) Current price of the stock is the PV of the expected dividends when discounted at the required return of 12% =
= 5/1.12+50/1.12^2 = $               44.32
b) Share price in 1 year's time = 50/1.12 = $               44.64
c) Dividends required per year = 44.32*5000*(0.12*1.12^2)/(1.12^2-1) = $   1,31,120.30
1st Year:
Dividends receivable = 5000*5 = $       25,000.00
Balance required = 131120.30-25000 = $   1,06,120.30
Number of shares to be sold at the end of the 1st year to get 106120.30 = 106120.30/44.64 = 2377.25 Shares
2nd Year:
Number of shares in hand = 5000-2377.25 = 2622.75 Shares
Dividend receivable = 2622.75*50 = $   1,31,137.65
Home made dividend is achieved by selling the required
number of shares to make the shortfall in the dividend
paid by the company.
Thus, in the 1st Year 2377.25 shares are sold and in the 2nd
Year the dividend receivable is equal to the liquidating
dividend on the remaining shares of 2622.75.
d) Dividend receivable at the end of the first year = $       25,000.00
Balance required for buying the house at EOY 1 = 220000-25000 = $   1,95,000.00
Number of shares to be sold to realize the balance amount = 195000/44.64 = 4368.28 Shares
e) Amount required to buy the house at EOY 2 = 220000+46800 = $   2,66,800.00
Amount of dividend receivable at EOY 2 = 2622.75*44.64 = $   1,31,137.65
Balance required $   1,35,662.35
This should be made up by reinvesting the first year total
dividends of $131120.30 for 1 year.
The return required for this investment = 135662.35/131120.30 -1 = 3.46%